# Linear Convergence of the Proximal Gradient Method for Composite Optimization Under the Polyak-Łojasiewicz Inequality and Its Variant

### v1

@@ 3.2 Convex Composite Case
In this section, we consider the performance of the PGM under Condition 2 and assume the PL inequality (3) holds. We still need the iteration format (9), the interpolation condition for closed, proper and convex functions (7), and the PL inequality (10) mentioned in Section 3.1. However, we replace (5) with the interpolation condition for $L$-smooth convex functions (6).
@@
Suppose that Condition 2 holds, $F$ satisfy the PL inequality (3) on $X=\{x:F(x)\leq F(x_{1})\}$, and $x_{2}$ is generated by Algorithm 1. Then the following holds.
@@
where $B_{i,j}$ is defined in (6), $C_{1,2}$ is defined in (7), $D_{2}$ is defined in (10), and the last inequality holds since $\gamma\in(0,\frac{3}{2L}]$. Because the terms after the nonnegative parameters $\alpha_{1},\alpha_{2},\beta_{1}$, and $\lambda_{1}$ are nonpositive, we have
@@
where $B_{i,j}$ is defined in (6), $C_{1,2}$ is defined in (7), $D_{2}$ is defined in (10), and the last inequality holds since $\gamma\in(\frac{3}{2L},\frac{2}{L})$. The remainder of the proof follows a similar approach to the previous case. ∎
@@
Comparison. Now let us compare the our convergence results with existing literature. Recently, Garrigos et al. [10] show that the convergence of the PGM (in our notation) is
@@
Therefore, we provide a tighter convergence rate for $\gamma\in(0,\frac{2}{L})$ than (11). In addition, the step size $\frac{1}{L}$ corresponds to the “optimal” step size for (11). When the step size is $\frac{1}{L}$, the contraction rate of (11) is $\frac{L}{L+\mu}.$ Proposition 4 shows that $\frac{3}{2L}$ give a rate $\frac{L}{L+3\mu}$, which is strictly better than (11). An illustration of the convergence rate computed by PEP with the bound (11) is shown in Figure 2.
@@ 4.1 Nonconvex Composite Case
$E_{i}:=(f_{i}+h_{i})-(f_{*}+h_{*})-\frac{1}{2\mu}\|\mathcal{G}_{i}\|^{2}\leq 0,i\in\{k,k+1\},$ (12)
where $\mathcal{G}_{i}=\mathcal{G}_{\gamma}(x_{i})$. Furthermore, plug $x_{k+1}=\mathrm{prox}_{\gamma h}(x_{k}-\gamma\nabla f(x_{k}))$ and iteration $x_{k+1}=x_{k}-\gamma(g_{k}+s_{k+1})$ into the proximal map-based residual $\mathcal{G}_{\gamma}(x_{k}):=\frac{1}{\gamma}\left(x_{k}-\mathrm{prox}_{\gamma h}(x_{k}-\gamma\nabla f(x_{k}))\right)$, we have
$\mathcal{G}_{\gamma}(x_{k})=\frac{1}{\gamma}(x_{k}-x_{k+1})=g_{k}+s_{k+1}.$ (13)
@@
Suppose that Condition 1 holds, $F$ satisfy the RPL inequality (4) on $X=\{x:F(x)\leq F(x_{1})\}$, and $x_{2}$ is generated by Algorithm 1. If $\gamma\in\left(0,\frac{2}{L}\right)$, then we have
@@
Let $\alpha_{1}=\beta_{1}=1,\lambda_{1}=\frac{\mu-\mu\left(L\gamma-1\right)^{2}}{L}$ and $\rho=\frac{L+\mu\left(L\gamma-1\right)^{2}-\mu}{L}$. By multiplying the inequalities by their respective factors and then summing them, we obtain that
@@
where $A_{1,2}$ is defined in (5), $C_{1,2}$ is defined in (7). Because the terms after the nonnegative parameters $\alpha_{1},\beta_{1},\lambda_{1}$ are nonpositive, we have the desired result. ∎
@@ 4.2 Convex Composite Case
In this section, we consider the performance of the PGM under Condition 2 and assume the RPL inequality (4) holds. Zhang and Zhang [18] proved a linear convergence rate for the PGM under the RPL inequality. By using PEP, we extend the step size interval to $(0,\frac{2}{L})$ and find a better convergence rate than that in [18]. We use the interpolation conditions of $\mathcal{F}_{0,\infty}$ in (7) and $\mathcal{F}_{0,L}$ in (6), respectively, and the discrete RPL inequality (12).
Lemma 1 (Theorem 3.5, [9]).
Let $F(x)=f(x)+h(x)$, where $f(x)\in\mathcal{F}_{0,L}$ and $h(x)\in\mathcal{F}_{0,\infty}$. For any $x\in\mathbb{R}^{d}$, it hold that $\|\mathcal{G}_{\gamma}(x)\|\leq d(0,\partial F(x)).$
Now we are ready to present our result.
@@
Suppose that Condition 2 holds, $F$ satisfy the RPL inequality (4) on $X=\{x:F(x)\leq F(x_{1})\}$, and $x_{2}$ is generated by Algorithm 1. Then the following holds.
@@
$\|g_{k+1}+s_{k+1}\|\geq d(0,\partial F(x_{k+1}))\geq\|\mathcal{G}_{k+1}\|$
from Lemma 1, and $\|\mathcal{G}_{k}\|^{2}=\|g_{k}+s_{k+1}\|^{2}$ in (13). We then have
@@
where $B_{1,2}$ is defined in (6), $C_{1,2}$ is defined in (7), $E_{i}$ is defined in (12) and the last inequality holds since $\gamma\in(0,\frac{1}{L}]$. Because the terms after the nonnegative parameters $\alpha_{1},\beta_{1},\lambda_{1}$ and $\lambda_{2}$ are nonpositive, we have
@@
A comparison of the convergence rate computed by PEP with the bound (15) is shown in Figure 2. The convergence rate in [18] is obtained using a refined descent lemma on the interval $(0,\frac{1}{L}]$, while we extend the interval to $(0,\frac{2}{L})$. Combined with Proposition 8, we find that by choosing a longer step size, one can achieve a better convergence rate than (15).
@@
From Theorems 3 and 7, we can observe that for convex composite functions, whether satisfying the PL or RPL inequality, the convergence rates are the same when the step size $\gamma\in(\frac{3}{2L},\frac{2}{L})$ but different in other intervals of step sizes. We remark that the main difference between the analysis of two cases are that (10) is used for the PL case and (14) is used for the RPL case.
@@ 5 Conclusion
In this paper, we provide explicit convergence rates of the PGM applied to composite functions using the PEP framework, focusing on two classes of the PL inequality. For nonconvex composite functions satisfying either the PL or RPL inequality, we present the first explicit linear convergence rate. Additionally, we derive tighter bounds on the linear convergence rate of PGM for minimizing convex composite functions under the PL or RPL inequality.

### v2

@@ 3.2 Convex Composite Case
We consider the performance of the PGM under Condition 2 and assume the PL inequality (3) holds. We still need the iteration format (9), the interpolation condition for closed, proper and convex functions (7), and the PL inequality (10) mentioned in Section 3.1. However, we replace (5) with the interpolation condition for $L$-smooth convex functions (6).
@@
Suppose that Condition 2 holds, $F$ satisfies the PL inequality (3) on $X=\{x:F(x)\leq F(x_{1})\}$, and $x_{2}$ is generated by Algorithm 1. Then the following holds.
@@
where $B_{1,2}$ and $B_{2,1}$ are defined in (6), $C_{1,2}$ is defined in (7), $D_{2}$ is defined in (10), and the last inequality holds since $\gamma\in(0,\frac{3}{2L}]$. Because the terms after the nonnegative parameters $\alpha_{1},\alpha_{2},\beta_{1}$, and $\lambda_{1}$ are nonpositive, we have
@@
where $B_{1,2}$ and $B_{2,1}$ are defined in (6), $C_{1,2}$ is defined in (7), $D_{2}$ is defined in (10), and the last inequality holds since $\gamma\in(\frac{3}{2L},\frac{2}{L})$. The remainder of the proof follows a similar approach to the previous case. ∎
@@
Comparison. Now let us compare our convergence results with existing literature. Recently, Garrigos et al. [11] show that the convergence of the PGM (in our notation) is
@@
Therefore, we provide a tighter convergence rate for $\gamma\in(0,\frac{2}{L})$ than (11). In addition, the step size $\frac{1}{L}$ corresponds to the “optimal” step size for (11). When the step size is $\frac{1}{L}$, the contraction rate of (11) is $\frac{L}{L+\mu}.$ Proposition 4 shows that $\frac{3}{2L}$ give a rate $\frac{L}{L+3\mu}$, which is strictly better than (11). An illustration of the convergence rate computed by PEP with the bound (11) is shown in Fig. 2.
@@ 4.1 Nonconvex Composite Case
$E_{i}:=f_{i}+h_{i}-F_{*}-\frac{1}{2\mu}\|\mathcal{G}_{i}\|^{2}\leq 0,$ (12)
where $\mathcal{G}_{i}=\mathcal{G}_{\gamma}(x_{i})$. Furthermore, plug $x_{i+1}=\mathrm{prox}_{\gamma h}(x_{i}-\gamma\nabla f(x_{i}))$ and iteration $x_{i+1}=x_{i}-\gamma(g_{i}+s_{i+1})$ into the proximal map-based residual $\mathcal{G}_{\gamma}(x_{i}):=\frac{1}{\gamma}\left(x_{i}-\mathrm{prox}_{\gamma h}(x_{i}-\gamma\nabla f(x_{i}))\right)$, we have
$\mathcal{G}_{\gamma}(x_{i})=\frac{1}{\gamma}(x_{i}-x_{i+1})=g_{i}+s_{i+1}.$ (13)
@@
Suppose that Condition 1 holds, $F$ satisfies the RPL inequality (4) on $X=\{x:F(x)\leq F(x_{1})\}$, and $x_{2}$ is generated by Algorithm 1. If $\gamma\in\left(0,\frac{2}{L}\right)$, then we have
@@
Let $\alpha_{1}=\beta_{1}=1$, and $\lambda_{1}=\frac{\mu-\mu\left(L\gamma-1\right)^{2}}{L}$. By multiplying the inequalities by their respective factors and then summing them, we obtain that
@@
where $A_{1,2}$ is defined in (5), $C_{1,2}$ is defined in (7), and $E^{\prime}_{1}$ is defined in (14). Because the terms after the nonnegative parameters $\alpha_{1},\beta_{1},\lambda_{1}$ are nonpositive, we have the desired result. ∎
@@ 4.2 Convex Composite Case
Next, we consider the performance of the PGM under Condition 2 and assume the RPL inequality (4) holds. Zhang and Zhang [23] proved a linear convergence rate for the PGM under the RPL inequality. By using PEP, we extend the step size interval to $(0,\frac{2}{L})$ and find a better convergence rate than that in [23]. We use the interpolation conditions of $\mathcal{F}_{0,\infty}(\mathbb{R}^{d})$ in (7) and $\mathcal{F}_{0,L}(\mathbb{R}^{d})$ in (6), respectively, and the discrete RPL inequality (12). Now we are ready to present our result.
@@
Suppose that Condition 2 holds, $F$ satisfies the RPL inequality (4) on $X=\{x:F(x)\leq F(x_{1})\}$, and $x_{2}$ is generated by Algorithm 1. Then the following holds.
@@
$\|g_{i+1}+s_{i+1}\|\geq d(0,\partial F(x_{i+1}))\geq\|\mathcal{G}_{i+1}\|$
from Lemma 1, and $\|\mathcal{G}_{i}\|^{2}=\|g_{i}+s_{i+1}\|^{2}$ in (13). We then have
@@
where $B_{1,2}$ is defined in (6), $C_{1,2}$ is defined in (7), $E_{1}$ and $E_{2}$ are defined in (12) and the last inequality holds since $\gamma\in(0,\frac{1}{L}]$. Because the terms after the nonnegative parameters $\alpha_{1},\beta_{1},\lambda_{1}$ and $\lambda_{2}$ are nonpositive, we have
@@
A comparison of the convergence rate computed by PEP with the bound (15) is shown in Fig. 2. The convergence rate in [23] is obtained using a refined descent lemma on the interval $(0,\frac{1}{L}]$, while we extend the interval to $(0,\frac{2}{L})$. Combined with Proposition 8, we find that by choosing a longer step size, one can achieve a better convergence rate than (15).
@@
From Theorems 3 and 7, we can observe that for convex composite functions, whether satisfying the PL or RPL inequality, the convergence rates are the same when the step size $\gamma\in(\frac{3}{2L},\frac{2}{L})$ but different in other intervals of step sizes. We remark that the main difference between the analysis of the two cases is that (10) is used for the PL case and (14) is used for the RPL case.
5 Numerical Experiments
In this section, we compare the performance of different step size strategies for the PGM through numerical experiments on various practical problems.
5.1 Numerical comparisons under the PL inequality
First, we consider a nonconvex composite problem, the sparse robust linear regression (SRLR) problem, which has been studied in prior work, such as Ochs et al. [16], Zhou et al. [24], and Khanh et al. [12]. The problem is formulated as:
$\min_{x\in\mathbb{R}^{d}}\sum_{k=1}^{n}\log\left((Ax-b)_{k}^{2}+1\right)+\lambda\|x\|_{1},$
where $A\in\mathbb{R}^{n\times d}$, $b\in\mathbb{R}^{n}$, $\lambda>0$ is the regularization parameter, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm. We generate $A\in\mathbb{R}^{n\times d}$ and $b\in\mathbb{R}^{n}$ with independent and identically distributed (i.i.d.) standard normal entries, where $n=200$ and $d=20$, and set the regularization parameter to $\lambda=0.1$. Fig. 4 shows the results of using PGM to solve the SRLR problem with different step sizes. Particularly, we report the suboptimality gap, $F(x_{i})-F(\hat{x})$, where $\hat{x}$ denotes the suboptimal solution obtained by running PGM for a large number of iterations with high tolerance.
As shown in Fig. 4, PGM initially exhibits sublinear convergence, and transitions to linear convergence as it approaches a limit point. Since the objective function is the sum of a logarithmic function composed with a polynomial and the $\ell_{1}$ norm, it possesses an o-minimal structure. According to Bolte et al. [3], functions with o-minimal structure satisfy the KL inequality. In particular, we verified that the Hessian of the smooth part of the objective is positive definite near the point of convergence, implying local strong convexity. This suggests that the KL exponent tends to $\frac{1}{2}$, confirming that the PL inequality holds in a neighborhood of the point of convergence. This observation explains the linear convergence behavior near the optimum. We further observe that the PGM with the “optimal” step size $\frac{\sqrt{3}}{L}$, as stated in Proposition 2, outperforms other step sizes.
Figure 3: Comparison on the SRLR problem.
Figure 4: Comparison on the LASSO problem.
Next, we consider a convex composite problem, the LASSO problem [21]:
$\min_{x\in\mathbb{R}^{d}}\frac{1}{2}\|Ax-b\|^{2}+\lambda\|x\|_{1},$
where $A\in\mathbb{R}^{n\times d}$, $b\in\mathbb{R}^{n}$, and $\lambda>0$. We generate $A$ and $b$ in the same way as in the SRLR problem and set $\lambda=0.1$. It is known that the objective function of the LASSO problem satisfies the KL inequality with exponent $\frac{1}{2}$ [13], which coincides with the PL inequality discussed in this paper. The results of using PGM with different step sizes are shown in Fig. 4. In particular, we report the optimality in the objective function, $F(x_{i})-F_{*}$, with $F_{*}$ being the optimal solution computed by the solver CVX. We observe that the PGM using the “optimal” step size $\frac{3}{2L}$, as given in Proposition 4, converges faster than other step sizes, including the widely used choice $\frac{1}{L}$. This illustrates the effectiveness of our theoretical step size in improving the convergence rate.
5.2 Numerical comparisons under the RPL inequality
Finally, we consider numerical comparisons under the RPL inequality. However, we have not yet identified meaningful practical nonconvex composite problems that satisfy the RPL inequality. As a result, we do not conduct numerical experiments for the nonconvex case. Instead, we focus on the convex setting and consider the elastic net problem introduced by Zou and Hastie [25]:
$\min_{x\in\mathbb{R}^{d}}\frac{1}{2}\|Ax-b\|^{2}+\lambda\|x\|_{1}+\frac{\delta}{2}\|x\|^{2},$
where $A\in\mathbb{R}^{n\times d}$, $b\in\mathbb{R}^{n}$, $\lambda>0$. We generate $A$ and $b$ in the same way as in the SRLR problem and set $\lambda=0.1$. This problem can be viewed as a combination of a smooth convex term $\frac{1}{2}\|Ax-b\|^{2}$ and a strongly convex regularizer $\lambda\|x\|_{1}+\frac{\delta}{2}\|x\|^{2}$. The objective function is strongly convex and thus satisfies the RPL inequality, and $\mu=\lambda_{\min}(A^{T}A)+\delta$.
Figure 5: Comparison on the elastic net problem ($\mu\leq L/9$).
Figure 6: Comparison on the elastic net problem ($L/9<\mu<L$).
We conduct experiments using the same setup as in the LASSO problem and compare the performance of PGM with different step sizes. Note that the “optimal” step size is either $\frac{3}{2L}$ if $\mu\leq\frac{L}{9}$, or $\frac{2}{\sqrt{L}(\sqrt{L}+\sqrt{\mu})}$ if $\frac{L}{9}<\mu<L$ in Proposition 8. We set $\delta=10^{-2}$ in Fig. 6 to correspond to the case where $\mu\leq\frac{L}{9}$, and $\delta=100$ in Fig. 6 to correspond to the case where $\frac{L}{9}<\mu<L$. In particular, we computed that the step size $\gamma\approx\frac{1.991}{L}$ is close to $\frac{2}{L}$ when $\delta=10^{-2}$, and $\gamma\approx\frac{1.404}{L}<\frac{3}{2L}$ when $\delta=100$. The results in both figures show that PGM with the “optimal” step size converges faster than with other step sizes, including the commonly used choice $\frac{1}{L}$.
@@ 6 Conclusion
In this paper, we provide explicit convergence rates of the PGM applied to composite functions using the PEP framework, focusing on two classes of the PL inequality. For nonconvex composite functions satisfying either the PL or RPL inequality, we present the first explicit linear convergence rate. Additionally, we derive tighter bounds on the linear convergence rate of PGM for minimizing convex composite functions under the PL or RPL inequality. Finally, we conduct numerical experiments to show that the theoretically “optimal” step sizes derived in this paper achieve better convergence rates than other step sizes.